# T2V-Turbo-v2: Enhancing Video Generation Model Post-Training through Data, Reward, and Conditional Guidance Design

### v2

@@ 2 Preliminary
Diffusion Sampling The sampling process from a latent video diffusion model (Ho et al., 2022b) can be treated as solving an empirical probability flow ordinary differential equation (PF-ODE) (Song et al., 2020b). The sampling process executes in a reversed-time order, starts from a standard Gaussian noise $\boldsymbol{z}_{T}$, and returns a clean latent $\boldsymbol{z}_{0}$.
$\mathrm{d}\boldsymbol{z}_{t}=\left[\boldsymbol{\mu}\left(t\right)\boldsymbol{z}_{t}+\frac{1}{2}\sigma(t)^{2}\boldsymbol{\epsilon}_{\psi}(\boldsymbol{z}_{t},\boldsymbol{c},t)\right]\mathrm{d}t,\quad\boldsymbol{z}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (1)
where $\boldsymbol{z}_{t}\sim p_{t}(\boldsymbol{z}_{t})$ is the noisy latent at timestep $t$, $\boldsymbol{c}$ is the text prompt, and $\boldsymbol{\mu}(\cdot)$ and $\sigma(\cdot)$ are the drift and diffusion coefficients, respectively. The denoising model $\boldsymbol{\epsilon}_{\psi}(\boldsymbol{z}_{t},\boldsymbol{c},t)$ is trained to approximate the score function $-\nabla\log p_{t}(\boldsymbol{z}_{t})$ via score matching, and is used to construct an ODE solver $\Psi$.
To improve the quality of conditional sampling, we can augment the denoising model $\boldsymbol{\epsilon}_{\psi}$ with classifier-free guidance (CFG) (Ho & Salimans, 2021) and the gradient of an energy function $\mathcal{G}$
$\hat{\epsilon_{\psi}}(\boldsymbol{z}_{t},\boldsymbol{c},t)=\epsilon_{\psi}\left(\boldsymbol{z}_{t},\boldsymbol{c},t\right)+\omega\left(\epsilon_{\psi}\left(\boldsymbol{z}_{t},\boldsymbol{c},t\right)-\epsilon_{\psi}\left(\boldsymbol{z}_{t},\varnothing,t\right)\right)+\lambda\nabla_{\boldsymbol{z}_{t}}\mathcal{G}\left(\boldsymbol{z}_{t},t\right),$ (2)
@@
Consistency Model (CM) (Song et al., 2023; Luo et al., 2023) is proposed to accelerate the sampling from a PF-ODE. Specifically, it learns a consistency function $\boldsymbol{f}:\left(\boldsymbol{z}_{t},\boldsymbol{c},t\right)\mapsto\boldsymbol{x}_{\epsilon}$ to directly map any $\boldsymbol{z}_{t}$ on the PF-ODE trajectory to its origin, where $\epsilon$ is a fixed small positive number. We can model $\boldsymbol{f}$ with a CM $\boldsymbol{f}_{\theta}$ and distill it from a pretrained diffusion model, e.g., a denoising model $\boldsymbol{\epsilon}_{\psi}$, by minimizing the consistency distillation (CD) loss.
Consider discretizing the time horizon into $N-1$ sub-intervals with $t_{1}=\epsilon<t_{2}<\ldots<t_{N}=T$, we can sample any $\boldsymbol{z}_{t_{n+k}}$ ($k$ is the skipping interval of the ODE solver $\Psi$) from the PF-ODE trajectory and obtain its solution with $\Psi$. Conventional methods (Luo et al., 2023; Wang et al., 2023b; Li et al., 2024a; b) augment $\Psi$ with CFG and the solution is given as below
$\hat{\boldsymbol{z}}_{t_{n}}^{\Psi,\omega}\leftarrow\boldsymbol{z}_{t_{n+k}}+(1+\omega)\Psi(\boldsymbol{z}_{t_{n+k}},t_{n+k},t_{n},\boldsymbol{c};\psi)-\omega\Psi(\boldsymbol{z}_{t_{n}+k},t_{n+k},t_{n},\varnothing;\psi).$ (3)
One can condition $\boldsymbol{f}_{\theta}$ on $\omega$, and formulate the CD loss by enforcing CM’s self-consistency property:
$L_{\text{CD}}\left(\theta,\theta^{-};\Psi\right)=\mathbb{E}_{\boldsymbol{z},\boldsymbol{c},\omega,n}\left[d\left(\boldsymbol{f}_{\theta}\left(\boldsymbol{z}_{t_{n+k}},\omega,\boldsymbol{c},t_{n+k}\right),\boldsymbol{f}_{\theta^{-}}\left(\hat{\boldsymbol{z}}_{t_{n}}^{\Psi,\omega},\omega,\boldsymbol{c},t_{n}\right)\right)\right],$ (4)
@@
Temporal Attention is introduced to model the temporal dynamic of a video. Given a batch of video latents $\boldsymbol{z}\in\mathbb{R}^{B\times F\times C\times H\times W}$ with batch size $B$, $F$ frames, $C$ channels, and spatial dimensions $H\times W$, the temporal attention operator first reshapes the features by merging the spatial dimensions into the batch size dimension, leading to $\bar{\boldsymbol{z}}\in\mathbb{R}^{(B\times H\times W)\times F\times C}$. Then, it performs self-attention (Vaswani et al., 2017) along the frame axis to derive the temporal attention as below
$\mathcal{A}(\boldsymbol{z})=\texttt{Attention}\left(Q\left(\bar{\boldsymbol{z}}\right),K\left(\bar{\boldsymbol{z}}\right),V\left(\bar{\boldsymbol{z}}\right)\right)\in\mathbb{R}^{(B\times H\times W)\times F\times F},$ (5)
where $Q$, $K$, and $V$ are the Query, Key, and Value heads. And $\mathcal{A}(\boldsymbol{z})$ satisfies $\sum_{j=1}^{F}\mathcal{A}(\boldsymbol{z})_{(p,i,j)}=1$.
MotionClone (Ling et al., 2024) represents the video motion as the temporal attention and enables the cloning of the motion from a reference video to control the video generation process. At its core, MotionClone augments the conventional classifier-free guidance (Ho & Salimans, 2021) with Primary temporal-attention guidance (PTA guidance) and Location-aware semantic guidance (LAS guidance). In this paper, we focus on the PTA guidance. Given a reference video, we first obtain its latent $\boldsymbol{z}_{t}^{\text{ref}}$ at the timestep $t$ via DDIM inversion (Song et al., 2020a). Then, we can derive the energy function $\mathcal{G}^{m}$ associated with PTA guidance at timestep $t$ as below
$\mathcal{G}^{m}\left(\boldsymbol{z}_{t},\boldsymbol{z}_{t}^{\text{ref}},t\right)=\left\|\mathcal{M}\circ\left(\mathcal{A}(\boldsymbol{z}_{t}^{\text{ref}})-\mathcal{A}(\boldsymbol{z}_{t})\right)\right\|_{2}^{2},$ (6)
where $\circ$ denotes the Hadamard product and $\mathcal{M}$ is the temporal mask set to mask everything except for the highest activation along the temporal axis of $\mathcal{A}(\boldsymbol{z}_{t}^{\text{ref}})$, i.e.,
$\mathcal{M}_{(p,i,j)}:=\begin{cases}1,&\text{ if }\mathcal{A}(\boldsymbol{z}_{t}^{\text{ref}})_{(p,i,j)}=\max_{k}\mathcal{A}(\boldsymbol{z}_{t}^{\text{ref}})_{(p,i,k)}\\ 0,&\text{ otherwise}\end{cases}$ (7)
@@ 3 Integrate Rewards and Conditional Guidance into Consistency Distillation
Figure 2: Training pipeline of our T2V-Turbo-v2. When augmenting the teacher PF-ODE solver $\Psi$ with CFG and motion guidance, we extract motion prior $\mathcal{A}(\boldsymbol{z}_{t_{n+k}}^{\text{ref}})$ from the training videos and distill it into the student CM $\boldsymbol{f}_{\theta}$ along with the CFG.
Conventional CD methods (Song et al., 2023; Luo et al., 2023; Li et al., 2024a; b) typically neglect the design space of conditional guidance methods and do not employ an energy function $\mathcal{G}$ to augment the ODE solver $\Psi$. With the conditional guidance from $\mathcal{G}$, we can derive the solution of the augmented ODE solver as $\hat{\boldsymbol{z}}_{t_{n}}^{\Psi,\omega,\lambda}\leftarrow\hat{\boldsymbol{z}}_{t_{n}}^{\Psi,\omega}+\lambda\nabla_{\boldsymbol{z}_{t}}\mathcal{G}\left(\boldsymbol{z}_{t},t\right)$ with $\lambda$ controlling the conditional guidance strength. And similarly, we further condition our CM $\boldsymbol{f}_{\theta}$ on $\lambda$, i.e., $\boldsymbol{f}_{\theta}:\left(\boldsymbol{z}_{\boldsymbol{t}},\omega,\lambda,\boldsymbol{c},t\right)\mapsto\boldsymbol{z}_{\mathbf{0}}$.
@@
$\hat{\boldsymbol{z}}_{t_{n}}^{\Psi,\omega,\lambda}\leftarrow\boldsymbol{z}_{t_{n+k}}&+(1+\omega)\Psi(\boldsymbol{z}_{t_{n+k}},t_{n+k},t_{n},\boldsymbol{c};\psi)-\omega\Psi(\boldsymbol{z}_{t_{n}+k},t_{n+k},t_{n},\varnothing;\psi)\\ &+\lambda\cdot\nabla_{\boldsymbol{z}_{t_{n+k}}}\mathcal{G}^{m}\left(\boldsymbol{z}_{t_{n+k}},\boldsymbol{z}_{t_{n+k}}^{\text{ref}},t_{n+k}\right)\end{aligned},$ (8)
@@ 3.1 Learning Objectives
Our preliminary experiment shows that the target network $\boldsymbol{f}_{\theta^{-}}$ can be removed without affecting the training stability. This is empirically significant, as it frees us from maintaining the EMA $\theta^{-}$, saving a substantial amount of GPU memory. With our $\boldsymbol{f}_{\theta}$ conditioned $\lambda$, our new CD loss can be derived by slightly modifying Eq. 4
$\mathcal{L}_{\text{CD}}\left(\theta;\Psi\right)=\mathbb{E}_{\boldsymbol{z},\boldsymbol{c},\omega,n}\left[d\left(\boldsymbol{f}_{\theta}\left(\boldsymbol{z}_{t_{n+k}},\omega,\lambda,\boldsymbol{c},t_{n+k}\right),\texttt{stop\_grad}\left(\boldsymbol{f}_{\theta}\left(\hat{\boldsymbol{z}}_{t_{n}}^{\Psi,\omega,\lambda},\omega,\lambda,\boldsymbol{c},t_{n}\right)\right)\right)\right],$ (9)
@@
$J\left(\theta\right)=\mathbb{E}_{\boldsymbol{z},\boldsymbol{c},n}\left[\beta_{\text{img}}\sum^{M}_{m=1}\mathcal{R}_{\text{img}}\left(\hat{\boldsymbol{x}}^{m}_{0},\boldsymbol{c}\right)+\beta_{\text{v}}\mathcal{R}_{\text{v}}\left(\hat{\boldsymbol{x}}_{0},\boldsymbol{c}\right)\right],\quad\hat{\boldsymbol{x}}_{0}=\mathcal{D}\left(\boldsymbol{f}_{\theta}\left(\boldsymbol{z}_{t_{n+k}},\omega,\lambda,\boldsymbol{c},t_{n+k}\right)\right),$ (10)
@@ 3.2 Data Processing and Training Procedures
It is important to note that calculating the gradient of an energy function $\mathcal{G}$ can be computationally expensive, consuming substantial memory. For example, using the motion guidance $\mathcal{G}^{m}$ as the energy function further requires performing DDIM inversion to obtain $\boldsymbol{z}_{t_{n+k}}^{\text{ref}}$, which is too expensive to be done during each training iteration. For example, MotionClone’s original implementation can consume over 40GB of GPU memory and require 3 minutes to perform the DDIM inversion.
Fortunately, we identify that the solution $\hat{\boldsymbol{z}}_{t_{n}}^{\Psi,\omega,\lambda}$ can be pre-calculated before training the CM $\boldsymbol{f}_{\theta}$. Appendix A describes the detailed procedures of our preprocessing procedures in Algorithm 1 and includes the pseudo-codes for training in Algorithm 2. Our training pipeline is depicted in Fig. 2.
@@ 4 Experimental Results
Settings. We distill our T2V-Turbo-v2 from VideoCrafter2 (Chen et al., 2024a). All our models are trained on 8 NVIDIA A100 GPUs for 8K gradient steps without gradient accumulation. We use a batch size of 3 to calculate the CD loss and 1 to optimize the reward objective on each GPU device. During optimization of the image-text reward model $\mathcal{R}_{\text{img}}$, we randomly sample 2 frames from each video by setting $M=2$. The learning rate is set to $51e-$ 5, and the guidance scale is defined within the range $[\omega_{\text{min}},\omega_{\text{max}}]=[5,15]$. We use DDIM (Song et al., 2020a) as our ODE solver $\Psi$, with a skipping step parameter of $k=5$. For motion guidance (MG), we set the motion guidance percentage $\tau=0.5$ and strength $\lambda=500$.

### v3

@@ 2 Preliminary
Diffusion Sampling The sampling process from a latent video diffusion model (Ho et al., 2022b) can be treated as solving an empirical probability flow ordinary differential equation (PF-ODE) (Song et al., 2020b). The sampling process executes in a reversed-time order, starts from a standard Gaussian noise $\bm{z}_{T}$, and returns a clean latent $\bm{z}_{0}$.
$\mathrm{d}\bm{z}_{t}=\left[\bm{\mu}\left(t\right)\bm{z}_{t}+\frac{1}{2}\sigma(t)^{2}\bm{\epsilon}_{\psi}(\bm{z}_{t},\bm{c},t)\right]\mathrm{d}t,\quad\bm{z}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (1)
where $\bm{z}_{t}\sim p_{t}(\bm{z}_{t})$ is the noisy latent at timestep $t$, $\bm{c}$ is the text prompt, and $\bm{\mu}(\cdot)$ and $\sigma(\cdot)$ are the drift and diffusion coefficients, respectively. The denoising model $\bm{\epsilon}_{\psi}(\bm{z}_{t},\bm{c},t)$ is trained to approximate the score function $\nabla\log p_{t}(\bm{z}_{t})$ via score matching, and is used to construct an ODE solver $\Psi$.
To improve the quality of conditional sampling, we can augment the denoising model $\bm{\epsilon}_{\psi}$ with classifier-free guidance (CFG) (Ho & Salimans, 2021) and the gradient of an energy function $\mathcal{G}$
$\hat{\epsilon}_{\psi}(\bm{z}_{t},\bm{c},t)=\epsilon_{\psi}\left(\bm{z}_{t},\bm{c},t\right)+\omega\left(\epsilon_{\psi}\left(\bm{z}_{t},\bm{c},t\right)-\epsilon_{\psi}\left(\bm{z}_{t},\varnothing,t\right)\right)+\lambda\nabla_{\bm{z}_{t}}\mathcal{G}\left(\bm{z}_{t},t\right),$ (2)
@@
Consistency Model (CM) (Song et al., 2023; Luo et al., 2023) is proposed to accelerate the sampling from a PF-ODE. Specifically, it learns a consistency function $\bm{f}:\left(\bm{z}_{t},\bm{c},t\right)\mapsto\bm{x}_{\epsilon}$ to directly map any $\bm{z}_{t}$ on the PF-ODE trajectory to its origin, where $\epsilon$ is a fixed small positive number. We can model $\bm{f}$ with a CM $\bm{f}_{\theta}$ and distill it from a pretrained diffusion model, e.g., a denoising model $\bm{\epsilon}_{\psi}$, by minimizing the consistency distillation (CD) loss.
Consider discretizing the time horizon into N1N-1 sub-intervals with $t_{1}=\epsilon<t_{2}<\ldots<t_{N}=T$, we can sample any $\bm{z}_{t_{n+k}}$ ($k$ is the skipping interval of the ODE solver $\Psi$) from the PF-ODE trajectory and obtain its solution with $\Psi$. Conventional methods (Luo et al., 2023; Wang et al., 2023b; Li et al., 2024a; b) augment $\Psi$ with CFG and the solution is given as below
$\hat{\bm{z}}_{t_{n}}^{\Psi,\omega}\leftarrow\bm{z}_{t_{n+k}}+(1+\omega)\Psi(\bm{z}_{t_{n+k}},t_{n+k},t_{n},\bm{c};\psi)-\omega\Psi(\bm{z}_{t_{n}+k},t_{n+k},t_{n},\varnothing;\psi).$ (3)
One can condition $\bm{f}_{\theta}$ on $\omega$, and formulate the CD loss by enforcing CM’s self-consistency property:
$L_{\text{CD}}\left(\theta,\theta^{-};\Psi\right)=\mathbb{E}_{\bm{z},\bm{c},\omega,n}\left[d\left(\bm{f}_{\theta}\left(\bm{z}_{t_{n+k}},\omega,\bm{c},t_{n+k}\right),\bm{f}_{\theta^{-}}\left(\hat{\bm{z}}_{t_{n}}^{\Psi,\omega},\omega,\bm{c},t_{n}\right)\right)\right],$ (4)
@@
Temporal Attention is introduced to model the temporal dynamic of a video. Given a batch of video latents $\bm{z}\in\mathbb{R}^{B\times F\times C\times H\times W}$ with batch size $B$, $F$ frames, $C$ channels, and spatial dimensions $H\times W$, the temporal attention operator first reshapes the features by merging the spatial dimensions into the batch size dimension, leading to $\bar{\bm{z}}\in\mathbb{R}^{(B\times H\times W)\times F\times C}$. Then, it performs self-attention (Vaswani et al., 2017) along the frame axis to derive the temporal attention as below
$\mathcal{A}(\bm{z})=\texttt{Attention}\left(Q\left(\bar{\bm{z}}\right),K\left(\bar{\bm{z}}\right),V\left(\bar{\bm{z}}\right)\right)\in\mathbb{R}^{(B\times H\times W)\times F\times F},$ (5)
where $Q$, $K$, and $V$ are the Query, Key, and Value heads. And $\mathcal{A}(\bm{z})$ satisfies $\sum_{j=1}^{F}\mathcal{A}(\bm{z})_{(p,i,j)}=1$.
MotionClone (Ling et al., 2024) represents the video motion as the temporal attention and enables the cloning of the motion from a reference video to control the video generation process. At its core, MotionClone augments the conventional classifier-free guidance (Ho & Salimans, 2021) with Primary temporal-attention guidance (PTA guidance) and Location-aware semantic guidance (LAS guidance). In this paper, we focus on the PTA guidance. Given a reference video, we first obtain its latent $\bm{z}_{t}^{\text{ref}}$ at the timestep $t$ via DDIM inversion (Song et al., 2020a). Then, we can derive the energy function $\mathcal{G}^{m}$ associated with PTA guidance at timestep $t$ as below
$\mathcal{G}^{m}\left(\bm{z}_{t},\bm{z}_{t}^{\text{ref}},t\right)=\left\|\mathcal{M}\circ\left(\mathcal{A}(\bm{z}_{t}^{\text{ref}})-\mathcal{A}(\bm{z}_{t})\right)\right\|_{2}^{2},$ (6)
where $\circ$ denotes the Hadamard product and $\mathcal{M}$ is the temporal mask set to mask everything except for the highest activation along the temporal axis of $\mathcal{A}(\bm{z}_{t}^{\text{ref}})$, i.e.,
$\mathcal{M}_{(p,i,j)}:=\begin{cases}1,&\text{ if }\mathcal{A}(\bm{z}_{t}^{\text{ref}})_{(p,i,j)}=\max_{k}\mathcal{A}(\bm{z}_{t}^{\text{ref}})_{(p,i,k)}\\ 0,&\text{ otherwise}\end{cases}$ (7)
@@ 3 Integrate Rewards and Conditional Guidance into Consistency Distillation
Figure 2: Training pipeline of our T2V-Turbo-v2. When augmenting the teacher PF-ODE solver $\Psi$ with CFG and motion guidance, we extract motion prior $\mathcal{A}(\bm{z}_{t_{n+k}}^{\text{ref}})$ from the training videos and distill it into the student CM $\bm{f}_{\theta}$ along with the CFG.
Conventional CD methods (Song et al., 2023; Luo et al., 2023; Li et al., 2024a; b) typically neglect the design space of conditional guidance methods and do not employ an energy function $\mathcal{G}$ to augment the ODE solver $\Psi$. With the conditional guidance from $\mathcal{G}$, we can derive the solution of the augmented ODE solver as $\hat{\bm{z}}_{t_{n}}^{\Psi,\omega,\lambda}\leftarrow\hat{\bm{z}}_{t_{n}}^{\Psi,\omega}+\lambda\nabla_{\bm{z}_{t}}\mathcal{G}\left(\bm{z}_{t},t\right)$ with $\lambda$ controlling the conditional guidance strength. And similarly, we further condition our CM $\bm{f}_{\theta}$ on $\lambda$, i.e., $\bm{f}_{\theta}:\left(\bm{z}_{\bm{t}},\omega,\lambda,\bm{c},t\right)\mapsto\bm{z}_{\mathbf{0}}$.
@@
$\begin{aligned} \hat{\bm{z}}_{t_{n}}^{\Psi,\omega,\lambda}\leftarrow\bm{z}_{t_{n+k}}&+(1+\omega)\Psi(\bm{z}_{t_{n+k}},t_{n+k},t_{n},\bm{c};\psi)-\omega\Psi(\bm{z}_{t_{n}+k},t_{n+k},t_{n},\varnothing;\psi)\\ &+\lambda\cdot\nabla_{\bm{z}_{t_{n+k}}}\mathcal{G}^{m}\left(\bm{z}_{t_{n+k}},\bm{z}_{t_{n+k}}^{\text{ref}},t_{n+k}\right)\end{aligned},$ (8)
@@ 3.1 Learning Objectives
Our preliminary experiment shows that the target network $\bm{f}_{\theta^{-}}$ can be removed without affecting the training stability. This is empirically significant, as it frees us from maintaining the EMA $\theta^{-}$, saving a substantial amount of GPU memory. With our $\bm{f}_{\theta}$ conditioned $\lambda$, our new CD loss can be derived by slightly modifying Eq. 4
$\mathcal{L}_{\text{CD}}\left(\theta;\Psi\right)=\mathbb{E}_{\bm{z},\bm{c},\omega,n}\left[d\left(\bm{f}_{\theta}\left(\bm{z}_{t_{n+k}},\omega,\lambda,\bm{c},t_{n+k}\right),\texttt{stop\_grad}\left(\bm{f}_{\theta}\left(\hat{\bm{z}}_{t_{n}}^{\Psi,\omega,\lambda},\omega,\lambda,\bm{c},t_{n}\right)\right)\right)\right],$ (9)
@@
$J\left(\theta\right)=\mathbb{E}_{\bm{z},\bm{c},n}\left[\beta_{\text{img}}\sum^{M}_{m=1}\mathcal{R}_{\text{img}}\left(\hat{\bm{x}}^{m}_{0},\bm{c}\right)+\beta_{\text{v}}\mathcal{R}_{\text{v}}\left(\hat{\bm{x}}_{0},\bm{c}\right)\right],\quad\hat{\bm{x}}_{0}=\mathcal{D}\left(\bm{f}_{\theta}\left(\bm{z}_{t_{n+k}},\omega,\lambda,\bm{c},t_{n+k}\right)\right),$ (10)
@@ 3.2 Data Processing and Training Procedures
It is important to note that calculating the gradient of an energy function $\mathcal{G}$ can be computationally expensive, consuming substantial memory. For example, using the motion guidance $\mathcal{G}^{m}$ as the energy function further requires performing DDIM inversion to obtain $\bm{z}_{t_{n+k}}^{\text{ref}}$, which is too expensive to be done during each training iteration. For example, MotionClone’s original implementation can consume over 40GB of GPU memory and require 3 minutes to perform the DDIM inversion.
Fortunately, we identify that the solution $\hat{\bm{z}}_{t_{n}}^{\Psi,\omega,\lambda}$ can be pre-calculated before training the CM $\bm{f}_{\theta}$. Appendix A describes the detailed procedures of our preprocessing procedures in Algorithm 1 and includes the pseudo-codes for training in Algorithm 2. Our training pipeline is depicted in Fig. 2.
@@ 4 Experimental Results
Settings. We distill our T2V-Turbo-v2 from VideoCrafter2 (Chen et al., 2024a). All our models are trained on 8 NVIDIA A100 GPUs for 8K gradient steps without gradient accumulation. We use a batch size of 3 to calculate the CD loss and 1 to optimize the reward objective on each GPU device. During optimization of the image-text reward model $\mathcal{R}_{\text{img}}$, we randomly sample 2 frames from each video by setting $M=2$. The learning rate is set to 1e51e-5, and the guidance scale is defined within the range $\omega_{\text{min}},\omega_{\text{max}}]=[5,15]$. We use DDIM (Song et al., 2020a) as our ODE solver $\Psi$, with a skipping step parameter of $k=5$. For motion guidance (MG), we set the motion guidance percentage $\tau=0.5$ and strength $\lambda=500$.